Determine the points of intersection using algebra.Check using graphing and substituition......a..3x=y-1=0,2x=y-6
Answer by checkley71(8403)

3X=Y-1
2X=Y-6 SUBTRACTING THE TWO EQUATIONS WE GET
------------
X=5 & 3*5=Y-1 OR 15=Y-1 OR Y=16 PROOF
2*5=16-6 OR 10=10